# Experimental Exploration: Investigating Cooperative Interaction Behavior Between Humans and Large Language Model Agents

### v1

@@ 7. Discussion
7.1. Human Cooperative Behavior Differences
The cooperation rate is the most significant metric to demonstrate the cooperative behavior situation. Our results suggest that humans prefer to make cooperative decisions with another human rather than rule-based AI agents or LLM agents in the competitive context. This finding is consistent with previous arguments that humans prefer to cooperate with other humans rather than AI (Xu et al., 2021; Li et al., 2022). The confirmation of Hypothesis 3(H3)indicates that gender also plays a significant role in human-AI cooperation, with important implications for both HCI and CSCW. Specifically, gender may influence users’ attitudes toward and strategies for cooperating with different purported-characteristic AI agents.
Understanding how gender influences interactions with AI agents offers valuable insights for developing social computing systems that better facilitate cooperation among diverse user groups. Future AI agent designs should incorporate these gender differences to enhance collaborative work. (Lim et al., 2023) mentioned to find out the key factors to improve AI-powered social computing systems. Our study supports that gender and purported characteristics should be considered. Designers could tailor AI agent behavior by adjusting communication styles and interaction approaches based on gender, such as modifying tone, assertiveness, or responsiveness to foster better cooperation (Ribino, 2023).
Additionally, the previous discussion of algorithm aversion mainly focuses on the context of AI systems acting as decision-making assistants for humans (Jussupow et al., 2020; Dietvorst et al., 2015). This study positioned AI agents as independent decision-making competitors to humans and found similar conclusions, which human had higher cooperation rate when cooperating with a purported human compared with a ruled-based or LLM-driven agent. The results of the semi-structured interviews indicate that participants believe the purported-human agent will exhibit certain value orientations. This finding is also corroborated by the work of Stapel and Koomen (2005), which verifies that people are more willing to cooperate with others who share similar goals. This provides us with a clear design implication: when designing AI agents, particularly in competitive cooperation contexts, it is an effective strategy to clearly communicate the agent’s goals to the participants to enhance cooperation efficiency.
For Figure 3, it is evident that during the first 15 rounds, participants exhibited a lower cooperation rate with the rule-based AI agent compared to the LLM agent. Structured interviews revealed the reasons underlying this phenomenon. When interacting with the rule-based AI agent, many participants relied on trial and error to discern its patterns (Young, 2009). In contrast, interactions with the LLM agent prompted a more cautious approach, as participants sought to establish stable cooperative relationships.
Another noteworthy observation is that participants were informed before the experiment that each set of games with the same opponent would last for 50 rounds. This knowledge led them to adopt more conservative and self-interested strategies, such as defection, particularly towards the end of the game—most notably after the 40th round. This strategic shift accounts for the marked decline in the cooperation rate observed in the final stages (Daskalakis et al., 2009). An important insight derived from this finding is that, in scenarios involving continuous competitive cooperation, situations with an uncertain endpoint tend to foster sustained cooperation better. This phenomenon aligns with existing research on game theory and iterated dilemmas, which often highlight the impact of finite versus indefinite time horizons on cooperative behavior (Pettit and Sugden, 1989). This area warrants further investigation to better understand the dynamics of cooperation in such contexts.
7.2. Proactive Favourable Behavior and Betrayal Forgiveness
Our findings reveal that the interaction between gender and the perceived characteristics of opponents significantly influences participants’ proactive, favorable behavior. Glikson and Woolley (2020) demonstrated that AI systems often attempt to exhibit friendly behavior to earn human trust and foster cooperation. Building on these insights, our results suggest that when LLM agents act as independent decision-makers in human-AI competitive cooperation scenarios, humans reciprocate by displaying favorable behaviors to enhance collaboration.
Participants were most inclined to exhibit the highest friendliness toward opponents they perceived as human. A comparison between purported human opponents and actual human opponents indicates that participants treated both groups equally in terms of friendliness, as both were perceived as genuine humans. Notably, participants maintained consistent levels of friendliness regardless of variations in cooperative behavior, even when the cooperation rate was low. Furthermore, our results demonstrate that participants favored the LLM agent more when paired with a purported rule-based AI agent. This preference stemmed from the perception that the LLM agent’s ability to analyze sentiment contributed to more favorable interactions (Xing, 2024).
For instance, designers can adopt an approach centered on emotional understanding (Vistorte et al., 2024). The structured interviews revealed that participants generally perceive LLM agents as possessing superior emotional sensing capabilities. Therefore, in practical applications, enabling LLM agents to demonstrate their willingness to assist and articulate their understanding of users may be effective strategies Bilquise et al. (2022). By emphasizing emotional understanding in the design process, LLM agents can more effectively perceive and respond to users’ emotional needs. Additionally, tailoring the expression styles and behavioral structures of AI agents based on users’ gender and their preferred types of AI collaboration not only enhances user satisfaction and trust but also promotes more effective human-computer collaboration Thomas et al. (2020). Consequently, this contributes to developing more intelligent, friendly, and empathetic AI agents. LLM agents should incorporate this mechanism to mitigate resentment, thereby improving human-AI cooperation.
7.3. Future and Limitation
This study offers valuable insights for AI agent designers and developers, emphasizing the importance of carefully considering target users, application contexts, and the purported characteristic of agents to meet practical requirements. The domain of AI-human collaborative interaction for task completion holds significant potential. Our research shifts the focus to competitive cooperation, addressing the growing prevalence of AI technologies capable of executing basic foundational tasks. While this represents a promising development, it also underscores the necessity of integrating more human-centered considerations into AI design. In future CSCW (Computer-Supported Cooperative Work) research, striking a balance between optimizing proclaimed characteristics and achieving cooperation efficiency remains an unresolved challenge. Although presenting an agent as a human can improve cooperative performance, it involves an element of deception, potentially leading to unforeseen risks if the agent’s true nature is disclosed. Additionally, designers may tailor an agent’s settings based on the gender of the human collaborator to enhance cooperation. However, it is crucial to ensure that such adjustments adhere to principles of transparency, fairness, and ethical responsibility.
However, our study is limited by the controlled experimental settings, specifically the Prisoners’ Dilemma game, and the particular LLM agents employed. These controlled conditions may not fully capture the complexity and variability of real-world cooperative interactions. Additionally, a learning effect is inevitable when participants play the game consecutively, potentially altering their original behavioral strategies. Another imperfect aspect of our experiment is the lack of effective approaches to optimize the design and measurement methods for better comparing human-human and human-AI performance in competitive cooperation scenarios, which presents a challenge.
To address these limitations, future studies should incorporate a broader range of LLM agents and more complex, realistic interaction scenarios. Based on this work, future research may explore the long-term effects of human interactions with LLM agents and examine the impact of continuous learning and adaptation in these agents. Given the potential widespread deployment of AI agents, policymakers should establish guidelines and ethical standards for their development and use. Ensuring that LLM agents’ behaviors and decision-making processes are reasonable is crucial for building public trust and acceptance.
8. Conclusions
With the continued proliferation of LLMs, these models are powered by an increasing number of applications and agents, driving a growing demand for deeper exploration of human-AI interaction. The transformation of AI from an assistant to an independent decision-making agent has already demonstrated significant potential and advantages in various contexts. It is crucial to investigate how humans are influenced by such agents, particularly in the domain of CSCW, where the dynamics of human-AI competitive cooperation merit further exploration.
This study successfully verified the three hypotheses proposed at the outset of our study. Moreover, our research has identified distinct behavioral patterns in human interactions with LLM agents in competitive cooperation scenarios. Specifically, we observed differences in participants’ cooperation rates, decision-making times, amicability, and acceptance of repair efforts after betrayal when interacting with different purported-characteristic agents. These findings contribute to advancing the understanding of human-AI interaction by illuminating the nuanced ways in which humans adapt their behaviors based on their perceptions of various agents.
It also addresses the relative lack of research on human-AI cooperation from a competitive perspective and provides experimental evidence of the value and potential of this line of inquiry. By uncovering the subtleties of human behavior in competitive cooperation, developers can design AI agents that promote improved collaboration and trust. More importantly, examining human cooperation with independent decision-making AI agents ahead of time offers critical insights and foresight for shaping ethical frameworks to guide future AI development.

### v2

@@ 5. Discussion
5.1. Affective Paradox in Human-AI Interaction: When LLM Agents Fail to ‘Get Me’
Our findings reveal a fundamental tension in human-LLM cooperation: participants exhibited asymmetric moral expectations toward LLM agents compared to rule-based AI agents. Participants initially exhibited higher cooperation rates with LLM agents than with rule-based counterparts; however, this trust deteriorated as the game progressed and LLM agents engaged in non-cooperative behavior. The sharp decline in cooperation after Round 15 (Fig.  6), coupled with intensified moral condemnation of LLM non-cooperation, suggests a moral ambiguity discount—where anthropomorphic projections initially enhance engagement but ultimately amplify distrust when agents exhibit behavior that violates expectations (Cohn et al., 2024).
Figure 6. Cooperation Rate by Declared Identity
The selection of the 4th-degree polynomial model is supported by its superior performance metrics, including the highest $R^{2}$ (.677) and $\text{Adj}_{R^{2}}$ (.668), as well as the lowest AIC (-288.122), BIC (-270.058), and $\text{CV}_{\text{RMSE}}$ (.090) (Vrieze, 2012).
This phenomenon extends the Computers as Social Actors (CASA) paradigm (Gambino et al., 2020): humans apply social heuristics to LLMs (e.g., reciprocity expectations in unsolicited cooperation acts), yet paradoxically penalize them for deviating from expected, predictable behavior. As P19 noted, “I wanted to cultivate a cooperative LLM”, participants anthropomorphized LLMs’ decision-making while also being keenly aware of their probabilistic nature—creating a cognitive tension (procedural transparency asymmetry) between human-like expectations and the machine’s underlying mechanics (Endsley, 2023). Interviews also revealed conflicting mental models: daily LLM users simultaneously viewed the technology as “creative partners” (P11) and “unaccountable black boxes” (P4), reflecting Dignum (2018) framework of moral uncertainty in AI ecosystems.
Notably, the anthropomorphism-complexity paradox emerged in questionnaire responses: 83.3% of participants attributed stronger moral condemnation to LLM agents for betrayal, citing their perceived “intentionality” (Q5). This aligns with participants’ strategic caution when interacting with LLMs (e.g., P30 avoiding early defection to avoid “labeling”). Such behaviors reflect an implicit negotiation of power dynamics, where humans attempt to "domesticate" LLMs through repeated interactions, mirroring patterns observed in human-to-human relationship-building in iterated games (Axelrod, 1981).
These findings underscore the duality of human-LLM collaboration: while anthropomorphism fosters engagement and cooperation, it also amplifies disillusionment when expectations are unmet. This tension highlights the need for more transparent and predictable AI behaviors to mitigate moral ambiguity and foster trust. Future research should explore design strategies to balance anthropomorphic engagement with procedural clarity, ensuring AI systems remain effective yet ethically accountable in cooperative contexts.
5.2. Fixed Rules vs. LLM in Systems: Preference Related to Anthropomorphic Features
Our questionnaire, as referenced in Section 4.3, posits that preferences for fixed-rule versus LLM-driven AI agents correlate with occupational norms and goal congruence. For instance, artists prioritize creative novelty and thus favor LLMs for generating unconventional ideas, whereas engineers value precision and reproducibility, aligning them with rule-based systems for analytical outputs. In addition, participants were most inclined to exhibit the highest friendliness toward opponents they perceived as human. This dichotomy extends to human-AI competitive cooperation dynamics. While prior research on algorithm aversion focuses on AI as decision-making assistants (Jussupow et al., 2020; Dietvorst et al., 2015), our results reveal that humans exhibit higher cooperation rates with purported human agents than with rule-based or LLM-driven counterparts. Semi-structured interviews suggest participants perceive human-like agents as sharing value orientations, a finding supported by Stapel and Koomen (2005), who demonstrate that cooperation thrives under goal similarity. A critical design implication emerges: Explicitly communicating AI agents’ objectives in competitive contexts enhances cooperation efficiency by mitigating perceived misalignment.
Besides, gender significantly modulates human-AI interaction strategies, offering actionable insights for socially adaptive AI systems. Confirmed in RQ2, gender differences manifest in users’ attitudes toward AI agents’ purported characteristics. These disparities reflect deeper psychological motivations—self-esteem, achievement, and social comparison (Wood, 1996; Suls et al., 2002)—which intensify as AI transitions from tool to societal participant. Furthermore, our results demonstrate that participants favored the LLM agent more when paired with a purported rule-based AI agent. This preference stemmed from the perception that the LLM agent’s ability to analyze sentiment contributed to more favorable interactions (Xing, 2024). To optimize collaboration, embedding motivational scaffolding into AI architectures could transform competition into productive synergy.
5.3. LLM Agents as Social Mirrors: Triggering Comparative Dynamics in Human Participants
Our results indicate that, as referenced in Section 4.3, participants were more likely to experience the self-enhancement effect when interacting with LLM agents as opposed to purported rule-based AI agents (Alicke and Sedikides, 2009). We hypothesize that this is due to participants’ stronger perception of LLM agents’ capabilities, leading them to view these agents as more autonomous decision-makers. This, in turn, induces a stronger social presence effect (Oh et al., 2018), wherein LLM agents prompt participants to invest more time and effort in reflection and decision-making, thereby enhancing the depth and caution of their thinking. This is clearly reflected in the analysis of decision latency.
In educational contexts, the integration of LLM agents can enhance student engagement and education performance (An et al., 2024; Song et al., 2024). As students interact with these agents, they may perceive them as more capable and intelligent than traditional educational tools, leading to greater investment in their learning. This heightened engagement can improve students’ ability to critically evaluate information and engage more deeply with content, much like the increased decision-making effort observed in our study when participants interacted with the LLM agent.
While the social presence effect can enhance engagement, there are potential risks of over-reliance on LLM agents in education (Hamdi, 2024; Rizvi, 2023). If students view the LLM as an infallible authority, it may hinder their ability to think independently and critically evaluate information. This dependency could reduce the development of essential problem-solving skills, as students may rely on the AI for answers instead of engaging in deeper cognitive processes. Additionally, the social presence effect might cause stress or anxiety if students feel pressured to match the LLM’s capabilities, potentially affecting their mental well-being and creating an unhealthy learning environment. The comparative dynamics of LLM agents may worsen disparities in learning. Students struggling with the cognitive demands may feel inadequate, while those excelling may become overly reliant on the agent, distancing themselves from genuine understanding. This could lead to unequal learning outcomes, undermining the potential benefits of the LLM.
Therefore, educators must strike a balance between the benefits of increased engagement and the need to foster independent thinking, ensuring that students learn to collaborate effectively with AI without becoming overly dependent on it. The goal is to help students develop the skills to solve complex problems while maintaining their ability to think independently.
5.4. Limitation and Future
While this study offers insights into human behavior in competitive cooperation with LLM agents, several limitations should be acknowledged.
First, the study was conducted in a controlled experimental environment using a repeated Prisoner’s Dilemma task. While this allowed for systematic observation of decision-making patterns, such settings may not fully reflect the richness and unpredictability of real-world human-AI collaboration. Additionally, repeated interactions introduced potential learning effects, which may have influenced participants’ behavioral strategies over time.
Second, the study focused on a limited set of agent types, with pre-defined “purported characteristics” such as being rule-based, human-like, or LLM-based. While these categories helped isolate specific effects, they may not capture the full diversity of AI systems or their evolving capabilities. Future studies could benefit from including a broader range of agent behaviors, learning capacities, and dynamic adaptation strategies to better approximate real-world deployment scenarios.
Another challenge lies in the current lack of robust comparative frameworks for assessing human-human and human-AI cooperation. The development of more refined experimental designs and metrics—such as trust calibration, adaptive strategy shifts, and transparency evaluation—remains an open direction for future research.
Furthermore, while our findings show that manipulating agent characteristics (e.g., by anthropomorphizing them or tailoring design based on gender) can influence cooperation outcomes, these approaches raise important ethical concerns. For example, portraying an agent as human may improve collaboration but introduces an element of deception, which could undermine user trust if revealed. Similarly, agent customization based on user demographics must be carefully examined to ensure fairness, transparency, and accountability.
One promising avenue emerging from this study is the analysis of reflection time as a behavioral metric. Our results suggest that reflection time is sensitive to both the perceived identity of the agent and the participant’s own characteristics (e.g., gender). Longer reflection times were observed when participants believed they were competing with a human agent, and male participants deliberated longer when interacting with same-gender opponents. These findings align with prior literature on human complacency in AI-assisted decision-making and suggest that reflection time may serve as a useful proxy for cognitive engagement and trust calibration in human-AI interaction.
Semi-structured interviews further revealed that participants often relied on prior assumptions or stereotypes when interpreting an agent’s behavior, which in turn shaped their responses. These insights highlight the need for future work to explore the cognitive and affective mechanisms underpinning human-agent interaction, especially in mixed-motive or competitive scenarios.
In summary, while this study provides a foundation for understanding behavioral adaptation in competitive Human-AI cooperation, future work should explore more ecologically valid settings, refine evaluative frameworks, and consider the broader ethical implications of agent design and deployment.
6. Conclusion
As large language models (LLMs) evolve beyond assistive roles to take on autonomous decision-making, new challenges emerge for human-AI interaction, especially in contexts involving both collaboration and competition. This study addressed three research questions concerning human behavior in competitive cooperation with LLM agents. Our findings revealed that participants’ cooperation rates, decision latency, unsolicited cooperation, and trust restoration varied depending on the perceived characteristics of the AI agents and the human’s gender. These differences suggest that users adapt their strategies based on how they interpret an agent’s intent and autonomy.
By examining mixed-motive interactions, this study contributes to a growing body of work that moves beyond traditional assistive paradigms. It offers empirical insights into how humans adjust their behavior when facing agentic AI systems, highlighting the need for context-sensitive and ethically grounded design. These findings underscore the importance of anticipating user expectations and perceptions as AI systems increasingly participate in foundational tasks, offering a timely foundation for responsible AI development and deployment.